# The Hao-Ng isomorphism theorem for reduced crossed products

### v1

@@ 1. Introduction
In this paper, we prove the reduced Hao-Ng isomorphism problem in complete generality, establishing the commutation of the reduced crossed product functor with the Cuntz-Pimsner C*-algebra functor. We briefly describe the problem at hand, where more details can be found in Section 2. Given a C*-correspondence $X$ over a C*-algebra ${\mathcal{B}}$, the smallest gauge-equivariant C*-algebra ${\mathcal{O}}_{X}$ generated by $X$ and ${\mathcal{B}}$ is called the Cuntz-Pimsner algebra of $X$. The Cuntz-Pimsner construction can be largely understood at the level of the underlying $\mathrm{C}^{*}$-correspondence [katsura2007ideal, katsura2004c], yet it encompasses a wide-variety of naturally occurring $\mathrm{C}^{*}$-algebras, including all crossed products by ${\mathbb{Z}}$ and topological graph algebras [katsura2004c, muhly2005topological, pimsner1996class]. Given a locally compact Hausdorff group $G$, a generalized gauge action on $X$ is an action $\alpha:G\curvearrowright{\mathcal{O}}_{X}$ that preserves the copies of $X$ and ${\mathcal{B}}$ inside ${\mathcal{O}}_{X}$, and allows us to form the reduced crossed product $\mathrm{C}^{*}$-correspondence $X\rtimes_{r,\alpha}G$ in a natural way. The reduced Hao-Ng isomorphism then asks whether the Cuntz-Pimsner algebra ${\mathcal{O}}_{X\rtimes_{r,\alpha}G}$ of $X\rtimes_{r,\alpha}G$ is isomorphic to the corresponding reduced crossed product ${\mathcal{O}}_{X}\rtimes_{r,\alpha}G$ of the Cuntz-Pimsner algebra of $X$.
@@
This tactic of proving commutation results for the C*-envelope was also witnessed by the first-named author together with Geffen and Eilers in [dor2020classification], where it was used to establish a connection between the seemingly distinct classification theories for C*-algebras and non-self-adjoint operator algebras. These types of links between the self-adjoint and non-self-adjoint theories have led to significant structure results in C*-algebra theory, including gauge-coaction co-univeraslity theorems for a variety of C*-algebras [dor2020tensor, dor2022c, dor2023normal, kakariadis2023couniversality, sehnem2022c], as well as various generalizations of Hao-Ng isomorphism theorems in the context of product systems [dor2020tensor, dor2022c, katsoulis2020product, kakariadis2024fock, li2022zappa].
@@ 2.2. C*-correspondences, Cuntz-Pimsner algebras, and their dynamics
Let ${\mathcal{B}},{\mathcal{C}}$ be $\mathrm{C}^{*}$-algebras and $X$ be a $\mathrm{C}^{*}$-correspondence over ${\mathcal{B}}$. A representation of $X$ is a pair $(\rho,t)$ consisting of a *-homomorphism $\rho:{\mathcal{B}}\rightarrow{\mathcal{C}}$ and a completely contractive linear map $t:X\rightarrow{\mathcal{C}}$ with the property that
@@ 4. Reduced crossed products and their C*-envelopes
Although the strategy presented in [katsoulis2021non] is shown here to be successful in resolving the reduced Hao-Ng isomorphsm in complete generality, the corresponding full Hao-Ng isomorphism is more difficult. Indeed, it is known that the full Hao-Ng isomorphism is equivalent to the commutation of the $\mathrm{C}^{*}$-envelope and the full crossed product for tensor algebras [katsoulis2021non, Theorem 4.9]. So, as before, it would suffice to prove an analogue of Theorem 4.3 for full crossed products. Unfortunately, this fails for arbitrary operator algebras [harris2019crossed, Theorem 5.7], while it is unknown whether it is valid for the class of tensor algebras. So, a more nuanced approach to the full Hao-Ng isomorphism would be necessary.
One such approach has been proposed for the subclass of so-called hyperrigid $\mathrm{C}^{*}$-correspondences. We say that a $\mathrm{C}^{*}$-correspondence $X$ is hyperrigid if all non-degenerate representations of ${\mathcal{O}}_{X}$ have the unique extension property when restricted to ${\mathcal{T}}_{X}^{+}$. Although many $\mathrm{C}^{*}$-correspondences are hyperrigid, as can be witnessed by Kim’s criterion [kim2021hyperrigidity], there are many interesting examples of C*-correspondences that are not hyperrigid. In fact, in a recent paper by Bilich [bilich2025maximality], several examples of C*-correspondences are shown to fail to satisfy Arveson’s hyperrigidity conjecture, and are hence not hyperrigid. For hyperrigid $\mathrm{C}^{*}$-correspondences, however, a notable reduction in establishing the full Hao-Ng isomorphism is to establish the coincidence of all full crossed products of ${\mathcal{T}}_{X}^{+}$ relative to admissible representations [katsoulis2021non, Theorem 4.9]. Thus, to establish the Hao-Ng isomorphism for full crossed products arising from hyperrigid $\mathrm{C}^{*}$-correspondences, it suffices to show that the full crossed product construction by a generalized gauge action, considered as an action on ${\mathcal{T}}_{X}^{+}$, does not depend on the choice of $\alpha$-admissible representation. Unfortunately, this reformulation has only been verified for row-finite graph $\mathrm{C}^{*}$-correspondences [katsoulis2021non, Corollary 5.3]. Thus, an important next step is to show that all full crossed products relative to admissible representations coincide for a larger class of tensor algebras, even for those arising from hyperrigid $\mathrm{C}^{*}$-correspondences.
Acknowledgments. The authors are grateful to Boyu Li for several fruitful discussions held at New Mexicao State University and electronically over Zoom, and to Boris Bilich for suggestions and comments on the paper. The authors are especially grateful to Raphael Clouâtre for pointing out a missing step in a preliminary version of the paper.

### v2

@@ 1. Introduction
In this paper, we resolve the reduced Hao-Ng isomorphism problem in complete generality, establishing the commutation of the reduced crossed product functor with the Cuntz-Pimsner C*-algebra functor. We briefly describe the problem at hand, where more details can be found in Section 2. Given a C*-correspondence $X$ over a C*-algebra ${\mathcal{B}}$, the smallest gauge-equivariant C*-algebra ${\mathcal{O}}_{X}$ generated by $X$ and ${\mathcal{B}}$ is called the Cuntz-Pimsner algebra of $X$. The Cuntz-Pimsner construction can be largely understood at the level of the underlying $\mathrm{C}^{*}$-correspondence [katsura2007ideal, katsura2004c], yet it encompasses a wide-variety of naturally occurring $\mathrm{C}^{*}$-algebras, including all crossed products by ${\mathbb{Z}}$ and topological graph algebras [katsura2004c, muhly2005topological, pimsner1996class]. Given a locally compact Hausdorff group $G$, a generalized gauge action on $X$ is an action $\alpha:G\curvearrowright{\mathcal{O}}_{X}$ that preserves the copies of $X$ and ${\mathcal{B}}$ inside ${\mathcal{O}}_{X}$, and allows us to form the reduced crossed product $\mathrm{C}^{*}$-correspondence $X\rtimes_{r,\alpha}G$ in a natural way. The reduced Hao-Ng isomorphism then asks whether the Cuntz-Pimsner algebra ${\mathcal{O}}_{X\rtimes_{r,\alpha}G}$ of $X\rtimes_{r,\alpha}G$ is isomorphic to the corresponding reduced crossed product ${\mathcal{O}}_{X}\rtimes_{r,\alpha}G$ of the Cuntz-Pimsner algebra of $X$.
@@
This tactic of proving commutation results for the C*-envelope was also witnessed by the first-named author together with Geffen and Eilers in [dor2020classification], where it was used to establish a connection between the seemingly distinct classification theories for C*-algebras and non-self-adjoint operator algebras. These types of links between the self-adjoint and non-self-adjoint theories have led to significant structure results in C*-algebra theory, including gauge-coaction co-universality theorems for a variety of C*-algebras [dor2020tensor, dor2022c, dor2023normal, kakariadis2023couniversality, sehnem2022c], as well as various generalizations of Hao-Ng isomorphism theorems in the context of product systems [dor2020tensor, dor2022c, katsoulis2020product, kakariadis2024fock, li2022zappa].
@@ 2.2. C*-correspondences, Cuntz-Pimsner algebras, and their dynamics
Let ${\mathcal{B}},{\mathcal{C}}$ be $\mathrm{C}^{*}$-algebras and $X$ be a $\mathrm{C}^{*}$-correspondence over ${\mathcal{B}}$. A representation of $X$ is a pair $(\rho,t)$ consisting of a non-degenerate *-homomorphism $\rho:{\mathcal{B}}\rightarrow{\mathcal{C}}$ and a completely contractive linear map $t:X\rightarrow{\mathcal{C}}$ with the property that
@@ 4. Reduced crossed products and their C*-envelopes
Although the strategy presented in [katsoulis2021non] is shown here to be successful in resolving the reduced Hao-Ng isomorphsm in complete generality, the corresponding full Hao-Ng isomorphism is more difficult. Indeed, it is known that the full Hao-Ng isomorphism is equivalent to the commutation of the $\mathrm{C}^{*}$-envelope and the full crossed product for tensor algebras [katsoulis2021non, Theorem 4.9]. As before, it would suffice to prove an analogue of Theorem 4.3 for full crossed products. Unfortunately, this fails for arbitrary operator algebras [harris2019crossed, Theorem 5.7], while it is unknown whether it is valid for the class of tensor algebras. Thus, a more nuanced approach would be necessary to resolve the full Hao-Ng isomorphism in complete generality.
Acknowledgments. The authors are grateful to Boyu Li for several fruitful discussions held at New Mexico State University and electronically over Zoom, and to Boris Bilich and Jamie Gabe for suggestions and comments on the paper. The authors are especially grateful to Raphaël Clouâtre for pointing out a missing argument in a preliminary version of the paper.